NAOJ GW Elog Logbook 3.2

Yuefan and Yuhang
Recently, we always have a problem of filter cavity green transmission power drop. It dropped from 5000 to 1500 recently. We figured out this problem today although we are always checking with the same method. The reason for GRtra power drop is too high gain.
After reducing the gain of filter cavity control loop while looking at filter cavity green transmission, we found the indicator of GRtra goes to 5000.
The incident power to filter cavity is 12.5mW, setting of the filter cavity loop is attached in figure 1 and 2(attenuation:1.68, gain:2.28). The locking bandwidth is 21kHz.

Yuhang and Yuefan
Since now we have the reflected beam coming from the filter cavity, it has astigmatism of ratio almost 2 between two axes. And also, it was jittering while testing.
It is reasonable that the range is larger at yaw direction while smaller at pitch direction due to astigmatism.
Anyway, the galvo works well with some oscillation. I attach here a video. https://drive.google.com/open?id=1sxK1SRzyOSfxDVE6_rPHDkVkjddDbnPS

While working remotely on loops optimization, I noticed that the BS oplev high frequency spectrum suddenly increased. See attached picture.
After some time (about ~20 min) it went back to the previous value. I suspect it might be due to an increase of environmental vibration maybe bought by the activation of some device.
Unfortunately I was not there to check. To be monitored.

I spent some time to optimize the new damp loops, whitout DC control, to be used when we want to engage AA or dithering control.
In the attached file the spectra with open and close loops for BS, INPUT and END mirror are reported together with the poles/zeros of the filters.
In some cases, an optimization (espacially to reduce overshoot) is still necessary.

To separate devices we put beneath the in-air bench, Matteo prepared a new shelf, we replaced it. The situation is now like the attached figure.

Yuhang and Yuefan
We tested the galvo for AA system.
We put a mirror after Faraday because there is still gate valve inside the vacuum tube. In this case(see attached figure 1), we could get a stable reflection. As mentioned by yuefan in the last entry, we had a problem of astigmatism in the beginning. We removed the second lens(-50mm lens), and then astigmatism disappeared.
Couldn't lock situation: If we found the light ERROR X or ERROR Y becomes light on(attached figure 2 and 3), it means we have a large offset in X or Y direction. So we couldn't lock loop and center beam.
Could lock situation: We found if error x is within 290mV and error y is within 730mV(attached figure 4 and 5), we could lock loop and center beam. But we found oscillation of around 30Hz or 1Hz after lock(attached figure 6 and 7).
The moment of lock is attached in figure 8.
Next step:
1. Get the real jittering beam and see if we could center beam.
2. Test galvo even with astigmatism.

Yuefan and Yuhang
Last Friday, for better understanding to beam situation, we decided to check again the simulation.
According to the beam size measured by Yuhang and Aritomi-san a while ago after the 500mm lens, we could get the original beam waist size and position of the reflection green beam. The beam waist is 2.5m before the 500mm lens, and the beam waist size is 440um which is different from what we expected (~1mm) and mearsured very long time ago by Yuhang and Marc(1008um).
Anyway then we checked the telescope I already designed with this beam, the result didn't seem good. The quadrant position for gouy phase 0 and 90 degree didn't change a lot, but the size became 100um on NF quadrant and 2.5mm on FF, both are bad for our requirement.
To increase the beam size on the NF quadrant, we could change the 500mm to longer one, but then this will limited mainly by the space we have. We decided to check again the beam profile next week when we get the cavity back to decide if we need to change the design or not.
Another thing we have to check is the astigmatism of the beam, because during the installation we found that the second lens which has a focal length of 50mm could induce a lot of astigmatism on the FF quadrant even we could not see by eyes the beam is off center. And if the beam already has a very strong astigmatism as measured in entry 1598, the final beam on the FF quadrant will have a really ugly shape. So after we check the beam again next, if it is still the same, we would like to check the injection path to see where the astigmatism come from.

[Aritomi, Yuhang, Yuefan, Eleonora, Matteo, Takahashi]
We opened PR chamber and measured BAB power in PR chamber. BAB power was fluctuating a bit and BAB power after dichoic mirror seems not reasonable, but anyway loss of dichroic mirror should be small. Loss of faraday in PR chamber is around 5% and other loss is less than 1%.
position | BAB power (uW) |
before viewport | 127 |
after viewport | 126.7 |
after faraday | 120.3 |
after dichroic mirror | 121 |
Then we put mirror after dichroic mirror and reflected BAB into faraday and maximized the reflection. BAB injection inside PR chamber is 131uW and BAB reflection is 116uW. The round trip loss is 11.5%.

Matteo, Simon
Yesterday, we finished to restart the translation stage after the power outage the day before.
After making sure that everything is working properly, we started to reconfigure the PCI bench for the polarization measurements:
- placing ND-filters in the beam-path before the shutter (-> order ~ 2)
- placing a Vantablack beam-dump at the same area to cover the ND-filter reflection
- putting a PBS after the first lens of the telescope
- putting a QWP and a HWP after the PBS
- placing a razor beam-dump before the second lens of the telescope to cover a ghost-beam coming from the ND-filters
- putting a 1" lens (f=150mm) together with a beam-tube on the detector stage
- putting a second PBS so that the beam-focus is ~ 2 cm behind the PBS
- placing a beam-dumb behind the blind-side of the PBS to cover back-scattering
- putting each a PSD (position sensitive detector) 2 cm behind the S-pol and P-pol side of the PBS
- Connecting the PSDs to the Lock-Ins and fine-tune the position of the PSDs with an oscilloscope (use the sum-output for the Lock-Ins and the X/Y output for the fine-tuning; beam should be centered)
After the setup is done, we calibrated the QWP and HWP in an iterative manner to produce a maximum in P-pol.
The data on the calibration are as follows:
S-pol PSD maximum voltage: 285 mV
S-pol PSD minimum voltage: 0.77 mV
P-pol PSD maximum voltage: 349 mV
P-pol PSD minimum voltage: 0.2 mV
That is comparable to what we got before.
(Note, that these number might be imrovable as we didn't spent so much time in adjusting the PBS position.)
We started a map of the Shinkosha TAMA#2 sample during the night. The results can be seen from the attached figures.
The polarization is relatively homogeneous. However, there are two main areas with different polarization-angles on two sides which leaves the center of the sample with the largest gradient in the polarization angle.
It is noteworthy that the distribution of polarization angles remembers a little bit on the 3-fold splitting visible in the absorption maps (see the elog entry 1633).

[Takahashi, Yuefan, Yuhang, Eleonora]
This morning with the help of Takahashi-san we closed PR chamber and restarted the vacuum. We will wait the vacuum to reach 10-6 torr (at BS vacuum monitor) to open the main gatevalve between BS and INPUT.
I upload some pictures taken to PR chamber.

Matteo, Simon
Attached to this report are the maps and distribution histograms of the absorption coefficient from the Shinkosha TAMA-sized TM #2.
As can be seen, we have now 7 different XY maps throughout the entire bulk. The YZ-map has a larger expansion in Y.
Although inhomogeneously distributed, the absorption is relatively low (~ 30 ppm/cm in the mean with some peaks exceeding 50 ppm/cm).
Please note that the colorscale is the same for all maps!

Yuefan and Yuhang
This morning we installed everything for the AA telescope before opening the chamber.
The final beam height is 225mm from the bench. All the optics position are only followed the design, we didn't have time to connect all the cables to check any signal. The second lens is installed on the rail, galvo and quadrants are all installed on the height changable posts.
The FF path is well aligned, the beam is hitting in the center of the quadrant. But the NF path, we still need to adjust a bit the mirrors of the galvo, to make the output beam on the same height and without any angle. It was not very easy to do because the position of the galvo is almost in the center of the bench.
The detail pictures are in the attachment. I found it was difficult to take a picture that shows all the components....I should also have taken some pictures from the other side of the bench, I will put them later...

[Takahashi, Aritomi, Yuefan, Matteo,Yuhang, Eleonora]
Today with the help of Takahashi-san we opened the BS and PR chambers.
First goal was to check a possible touching of the intermediate mass.
PR chamber
We found the intermediate mass was almost touching the magnet cage. Takahashi-san adjusted the cage position to leave more space between the cage and the intermediate mass. From the spectra taken before and after there is no evidence of improvement (pic1), but the spectra after the change have been taken with the chamber still open. So we will repet the measument in vacum to confirm the results.
BS chamber
BS intermediate mass seems fine.
The second activity was the assesment of losses inside PR chamber, from viewport, Faraday, dichroic mirror etc. Aritomi and Yuhang will report about that.
We closed the BS chamber and left the PR one open with the cleanboth. We wll close it tomorrow moring.

Yestarday afternoon at ~17.15 we had a blackout in TAMA due to a severe lightning storm. The power came back after ~10 minutes.
We spent the rest of the afternoon recovering the system. For the filter cavity everything seems fine. Takahashi-san came to help with vacuum system recovery.

I re-tried lock, but could not.
At first, I adjusted the offset of servo to set the monitored error signal to zero.
Also I minimize the gain of servo.
I tuned the laser frequency around the resonance, and turned on the integrator.
I tried several times with different conditions (inverted or not, and so on..), however, could not lock...
There may be some problems such as high finess, large gain, etc...
Actually, the efficiency of laser PZT is 50MHz/V, and the slope of error signal is 4.2e-7V/Hz.
Therefore, it needs to optimize the gain of the servo.

I found that green into OPO is completely misaligned. I recovered the alignment by measuring parametric amplification. Here is the information of parametric amplification of BAB.
green power (mW) | OPO temperature (kOhm) | p pol PLL (MHz) | Vmax (V) | Vmin (V) | parametric gain |
0 | 7.2 | 305 | 0.112 | 0.112 | 1 |
40 | 7.185 | 165 | 1.82 | 0.04 | 16.3 |

I installed the lens (f=50mm), QWP, and mirror for another double-pass AOM.
They are roughly located to compose a cat's eye retroreflector, but not tuned yet.

I tried PDH lock with a servo borrowed from TAMA, but could not lock.
At that time I found the laser power was larger or smaller than usual.
I monitored output voltage from the servo, and it was about 14V which saturated PZT frequency tunig range.
I played with gain, but it did not affect the output.
Actually, I did not adjust the offset before turnig on the integrators.
So I will re-try the lock as Matteo suggested; tune the offset first, then turn on the integrators with smaller gain around the resonance.
I re-tried lock, but could not.
At first, I adjusted the offset of servo to set the monitored error signal to zero.
Also I minimize the gain of servo.
I tuned the laser frequency around the resonance, and turned on the integrator.
I tried several times with different conditions (inverted or not, and so on..), however, could not lock...
There may be some problems such as high finess, large gain, etc...
Actually, the efficiency of laser PZT is 50MHz/V, and the slope of error signal is 4.2e-7V/Hz.
Therefore, it needs to optimize the gain of the servo.

Simulink model and MEDM screen for AA is now completed and ready to be tested with quadrants.
Note: while restarting the model last sunday I was not able to restore the screeshot file taken just before the restart. I didn't find any problem with the command but some of the epics channels remained blank. I check the file and see that some of that channels ware actually missing. I don't know why it happend. Anyway I restored a previous snapshot file. I restarted the model and restored snapshot several times after that and this problem never showed up again.

As we decided to try the dithering technique to keep long term alignment of the cavity, in the past day I realized the software necessary for the job. Both simulink model and MEDM interface. I tried to reuse code from KAGRA as much as possible.
We have how the possibility to send 4 lines for pitch and 4 lines for yaw to each of the 4 suspended mirror. and we can demodulate a cavity signal (to be chosen between GREEN TRA, IR TRA and FC CORR) at the frequency of these lines. Then we can decide to which mirrors to send the correction after properly filtering it.
In order to implement the dithering we should switch from DC local controls for mirrors to AC local control. I quickly realized a preliminary, non optimized version of these AC loops (except PR), which I hope to improve soon. Pic 3.